# QDM: Quadtree-Based Region-Adaptive Sparse Diffusion Models for Efficient Image Super-Resolution

### v1

@@ 4 Experiments
Table 1: Quantitative comparisons on LSDIR-Test and RealSR datasets. Best and second best: bold and underline. The number next to each method (e.g., -15 or -50) represents the number of inference steps used in that method.
Method LSDIR-Test RealSR
@@
StableSR-50 [33] 20.40 0.5448 0.2630 0.6821 68.6041 26.27 0.7755 0.2671 0.5209 60.1758
DiffBIR-50 [19] 21.28 0.5222 0.2813 0.7406 72.8067 24.83 0.6642 0.3864 0.6857 65.3934
SeeSR-50 [38] 21.36 0.5506 0.2686 0.7223 73.2258 26.20 0.7555 0.2806 0.6824 66.3757
OSEDiff-1 [39] 21.17 0.5517 0.2723 0.6896 72.1574 23.89 0.7030 0.3288 0.7008 65.4806
InvSR-1 [42] 20.23 0.5413 0.2747 0.7213 73.1948 24.50 0.7262 0.2872 0.6918 67.4586
SinSR-1 [36] 21.99 0.5750 0.2520 0.6683 69.2179 26.02 0.7097 0.3993 0.6634 59.2981
@@
BSRGAN [45] 22.05 0.5795 0.2929 0.5898 69.1739 26.51 0.7746 0.2685 0.5439 63.5869
RealESRGAN [35] 21.54 0.5840 0.2744 0.5887 69.4185 25.85 0.7734 0.2728 0.4898 59.6803
SwinIR [17] 21.53 0.5901 0.2652 0.6096 69.0808 26.43 0.7864 0.2513 0.4635 59.6491
LDM-50 [25] 22.00 0.5700 0.2776 0.5791 66.1659 26.75 0.7711 0.2945 0.4907 54.3910
ResShift-15 [44] 22.21 0.5879 0.2617 0.6398 68.8254 26.81 0.7675 0.3418 0.5714 57.6431
ResShift-4 [43] 22.37 0.6006 0.2462 0.6403 67.5648 25.77 0.7453 0.3395 0.5994 57.5536
QDM-L-15 22.17 0.5973 0.2441 0.6418 69.3115 25.82 0.7434 0.2925 0.6763 60.7319
@@ 4.1 Experimental Setup
Datasets. We train our proposed model on ×4 real-world super-resolution (SR) tasks , ×4 and x8 medical image SR tasks. For the real-world SR task, the training data integrates six established benchmarks: LSDIR [16], DIV2K [1], DIV8K [9], OutdoorSceneTraining [34], Flicker2K [18], and a subset of 10,000 facial images from the FFHQ dataset [13].
For the medical CT SR task, we employ clinical CT scans from two well-established segmentation challenges: HaN-Seg [24] (focused on head-and-neck organ segmentation) and SegRap2023 [21] (targeting naso-pharyngeal cancer radiotherapy planning). All medical images are partitioned into training and test sets using a 9:1 ratio to ensure rigorous evaluation.
To comprehensively assess real-world performance, we utilize two dedicated test benchmarks: RealSR [3] which contains 100 authentic LR-HR pairs captured using Canon 5D3 and Nikon D810 DSLRs, and RealSet80 [44] which curates 80 challenging low-resolution images from diverse internet sources and existing datasets.
Data Processing. For the real-world SR task, we adopt fixed-size $512\!\times\!512$ image crops during training. Following recent work [41], training images exceeding this resolution undergo either: (1) direct random cropping, or (2) resizing the shorter edge to 512 pixels before random cropping. We synthesize LR/HR pairs using the degradation pipeline from Real-ESRGAN [35].
Beyond real-world test benchmarks, we construct LSDIR-Test using center-cropped $512\!\times\!512$ images from the LSDIR test set [16], processed through identical degradation parameters as the training data.
For the medical CT SR task, we use the first degradation stage of Real-ESRGAN pipeline [35] simulating clinical noise artifacts. Both training and test images are first downsampled to $512\!\times\!512$ resolution as HR references, then processed through the adopted pipeline to obtain the corresponding LR image. Thus, we can create two benchmarks: Med-SR4 (×$\times 4$ super-resolution task) and Med-SR8 (×$\times 8$ super-resolution task).
Evaluation Metrics. The performance of different methods was evaluated using both reference-based and non-reference metrics. For reference-based assessment, we employed PSNR, SSIM [37], and LPIPS [46]. To better align with human perceptual judgments in generative super-resolution, we also incorporated non-reference metrics: CLIPIQA [32] and MUSIQ [14].
Implementation Details. We provide the implementation details of training process in Appendix B.
@@
Benchmarks Methods Metrics MACs(T)
@@
Med-SR4 (128→512) LDM-50 [25] 34.09 0.9751 0.0111 10.89
ResShift-15 [44] 35.84 0.9828 0.0097 3.27
QDM-S-15 33.99 0.9766 0.0162 0.37
QDM-B-15 35.10 0.9804 0.0147 1.42
QDM-L-15 35.75 0.9824 0.0120 5.00
Med-SR8 (64→512) LDM-50 [25] 31.00 0.9635 0.0162 10.89
ResShift-15 [44] 32.63 0.9728 0.0156 3.27
QDM-S-15 30.39 0.9617 0.0240 0.34
QDM-B-15 31.69 0.9681 0.0198 1.29
QDM-L-15 33.05 0.9743 0.0159 4.54
@@ 4.2 Model Analysis
We benchmark our method against contemporary super-resolution approaches spanning two dominant paradigms: (1) Pretrained-Prior-Based Methods, including StableSR [33], DiffBIR [19], SeeSR [38], OSEDiff [39], InvSR [42] and SinSR [36], which inherit generative priors from Stable Diffusion [25, 23] or employ distillation from a pretrained model; (2) Training-from-scratch Methods, comprising BSRGAN [45], RealESRGAN [35], and SwinIR [17], LDM [25], and ResShift [44, 43]. For sampling consistency, we configure 50 diffusion steps for StableSR, DiffBIR, SeeSR, and LDM. Method-specific configurations are preserved for OSEDiff (1 steps), InvSR (1 steps), SinSR (1 step), and ResShift (15 steps) following their official implementations.
As demonstrated by the qualitative and quantitative results in Fig. 4, Table 1, and Table 2, our proposed model outperforms or achieves comparable performance to state-of-the-art training-from-scratch methods across multiple benchmarks. While it still falls short of prior-based methods, it is worth emphasizing that such methods face significant practical limitations in medical imaging applications due to the absence of foundational generative models (e.g., Stable Diffusion) tailored to this domain. Furthermore, Table 2 highlights the computational efficiency of our QDM-B/S model: it delivers competitive performance relative to LDM and ResShift while requiring substantially less computational overhead. In contrast, lightweight methods like BSRGAN and SwinIR, though computationally efficient, achieve markedly inferior results. We show more experimental results on real-world benchmarks in Appendix C.
@@ 4.3 Ablation Studies
Quadtree mask threshold. As detailed in Section 3.3, the quadtree mask density controls computation allocation by adaptively partitioning images into homogeneous regions (processed via efficient upstream) and detail-rich areas (handled by high-resolution downstream). Table 3 demonstrates this on Med-SR8: At $s=0$ (lossless mode), QDM-L matches full-mask performance while cutting computation by $40.3\%$, confirming non-uniform prediction suffices for medical SR. Increasing $s$ to $0.90$ reduces computation to $15.5\%$ baseline (6.5×6.5$\times$6.5 × less) with only $0.34$ dB PSNR drop, favoring efficiency-critical clinical use. To balance performance and efficiency, we set the threshold $s=0.15$ for all medical SR experiments.
@@
Threshold PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ %Density MACs(T)
Full Mask 33.06 0.9744 0.0144 100.00 18.32
0.00 33.06 0.9744 0.0144 49.14 10.94
0.15 33.05 0.9743 0.0159 9.59 4.54
0.30 33.01 0.9739 0.0174 7.64 4.09
0.60 32.85 0.9728 0.0201 4.56 3.65
0.90 32.67 0.9719 0.0221 2.25 2.84
@@
Benchmark Param PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ MACs(T)
Med-SR4 (128→512) Patch (4) 35.18 0.9806 0.0113 4.10
Patch (8) 35.10 0.9804 0.0115 3.14
Window (4) 34.91 0.9797 0.0119 2.99
Window (8) 35.10 0.9804 0.0115 3.14
Window (16) 35.15 0.9807 0.0114 3.58
Med-SR8 (64→512) Patch (4) 31.88 0.9690 0.0172 4.07
Patch (8) 31.69 0.9682 0.0178 3.11
Window (4) 31.55 0.9673 0.0181 2.94
Window (8) 31.69 0.9682 0.0178 3.11
Window (16) 31.73 0.9683 0.0175 3.56
@@ 5 Conclusion
In this work, we present QDM, a novel diffusion-based super-resolution (SR) method that leverages quadtree-based region-adaptive sparse diffusion to enhance detail-rich image regions while minimizing computational overhead in homogeneous areas. The proposed framework incorporates two key innovations: 1) a quadtree-based the sparse diffusion model that rapidly detects detail-rich regions and only perform diffusion process in such areas, and 2) a two-stream architecture that decouples global and local feature extraction. The upstream branch captures broad contextual information using large patch sizes, while the downstream branch employs cross-attention to refine high-detail chunks identified via quadtree masks. This design achieves two advantages. First, it eliminates the need to store high-dimensional feature maps, significantly reducing memory overhead. Second, it dynamically minimizes computation by selectively processing only salient regions in the downstream branch. Extensive experiments on real-world SR and medical CT reconstruction tasks demonstrate that QDM outperforms existing training-from-scratch diffusion models in both reconstruction quality and computational efficiency. These results underscore the potential of region-adaptive diffusion strategies for resource-conscious high-fidelity image enhancement.
Supplementary Material
@@ Appendix A Different Configurations of QDM
Table 5: Details of QDM models. Our configurations for Small (S), Base (B) and Large (L) models are based on those of the DiT model [22]. To achieve a comparable parameter count, we halve the number of layers from the original DiT configurations.
Model Layers $N$ Hidden size Heads #Parameters(M)
@@ Appendix B Implementation Details
In real-world SR task, following prior work (LDM [25], ResShift [44]), our architecture operates in latent space using a Vector Quantized GAN (VQGAN, [6]) with a downsampling factor of 4. The model was trained for 150,000 iterations using the dual-stream training objective (Eq. 6), with a batch size of 64 on eight NVIDIA A100 80GB PCIe GPUs. We employed the Adam optimizer with a learning rate of $5\times 10^{-5}$, followed by a 50,000-iteration fine-tuning phase using the perceptual training objective (Eq. 7). For all real-world experiments, we set the quadtree threshold $s=0.00$.
For medical CT scan super-resolution, we first trained a KL-regularized VAE  [15, 25] on the medical dataset with a 4$\times$ downsampling factor. Our proposed model was then trained in this latent space for 50,000 iterations. To ensure fair comparison, ResShift and LDM were reimplemented and trained under identical conditions (50k iterations), while BSRGAN [45] and SwinIR [17] were trained for 100,000 iterations to account for their distinct optimization requirements. For all medical SR experiments, we set the quadtree threshold $s=0.15$.
For both tasks, the upstream patch size and downstream window size were set to 8. The low-resolution (LR) image was concatenated to the latent feature map before being passed to QDM, with a maximum of 64 chunks processed in parallel. Loss weightings $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ were all set to 1.
Besides, Figure 6 provide a detailed structure of cross-attention block presented in Figure 2. Each cross-attention block processes features sequentially through three components: a local intra-window self-attention, a cross-attention mechanism that integrates global context from upstream features, and a feedforward network. It’s important to note that the upstream global features are pre-projected into reusable key-value (kv) pairs to minimize redundant computations. The MLP ratio is set to the default value of 4.

### v2

@@ 4 Experiments
Table 1: Quantitative comparisons on the LSDIR-Test, RealSR, and RealSet80 datasets. The best and second-best results within each paradigm are highlighted in bold and underlined, respectively. The number following each method name (e.g., -15 or -50) indicates the number of inference steps used. All results are averaged over three runs with different random seeds.
Method LSDIR-Test RealSR RealSet80
@@
StableSR-50 [wang2024exploiting] 20.40 0.5448 0.2630 0.6821 68.6041 0.5209 60.1758 0.6214 62.7613
DiffBIR-50 [lin2024diffbir] 21.28 0.5222 0.2813 0.7406 72.8067 0.6857 65.3934 0.7404 67.9806
SeeSR-50 [wu2024seesr] 21.36 0.5506 0.2686 0.7223 73.2258 0.6824 66.3757 0.7114 69.7658
OSEDiff-1 [wu2025one] 21.17 0.5517 0.2723 0.6896 72.1574 0.7008 65.4806 0.7093 68.8022
InvSR-1 [yue2024arbitrary] 20.23 0.5413 0.2747 0.7213 73.1948 0.6918 67.4586 0.7291 69.8055
SinSR-1 [wang2024sinsr] 21.99 0.5750 0.2520 0.6683 69.2179 0.6634 59.2981 0.7228 64.0573
@@
BSRGAN [zhang2021designing] 22.05 0.5795 0.2929 0.5898 69.1739 0.5439 63.5869 0.6263 66.6288
RealESRGAN [wang2021real] 21.54 0.5840 0.2744 0.5887 69.4185 0.4898 59.6803 0.6189 64.4957
SwinIR [liang2021swinir] 21.53 0.5901 0.2652 0.6096 69.0808 0.4635 59.6491 0.6050 64.6984
LDM-50 [rombach2022high] 22.00 0.5700 0.2776 0.5791 66.1659 0.4907 54.3910 0.5511 55.8246
ResShift-15 [yue2024resshift] 22.23 0.5889 0.2620 0.6370 68.7279 0.5998 58.6835 0.6693 62.8512
QDM-L-15 22.16 0.5958 0.2452 0.6444 69.1535 0.6983 61.0735 0.6918 62.2867
@@ 4.1 Experimental Setup
Datasets. We train our proposed model on ×4 real-world super-resolution (SR) tasks , and ×4 and x8 medical image SR tasks. For the real-world SR task, the training data integrates six established benchmarks: LSDIR [li2023lsdir], DIV2K [agustsson2017ntire], DIV8K [9021973], OutdoorSceneTraining [wang2018recovering], Flickr2K [Lim_2017_CVPR_Workshops], and a subset of 10,000 facial images from the FFHQ dataset [karras2017progressive].
For the medical CT SR task, we employ clinical CT scans from two well-established segmentation challenges: HaN-Seg [podobnik2023han] (focused on head-and-neck organ segmentation) and SegRap2023 [luo2025segrap2023] (targeting naso-pharyngeal cancer radiotherapy planning). All medical images are partitioned into training and test sets using a 19:1 ratio to ensure rigorous evaluation.
To comprehensively assess real-world performance, we utilize two dedicated test benchmarks: RealSR [cai2019toward] which contains 100 authentic LR-HR pairs captured using Canon 5D3 and Nikon D810 DSLRs, and RealSet80 [yue2024resshift] which curates 80 challenging low-resolution images from diverse internet sources and existing datasets. We also construct LSDIR-Test benchmark using center-cropped $512\!\times\!512$ images from the LSDIR test set [li2023lsdir]. For the medical CT SR task, we create two benchmarks: Med-SR4 ($\times 4$ super-resolution task) and Med-SR8 ($\times 8$ super-resolution task). More details about data preprocessing, evaluation metrics as well as training process are in Appendix B.
@@
Benchmarks Methods Metrics FLOPs(T)
@@
Med-SR4 (128→512) LDM-50 [rombach2022high] 34.09 0.9751 0.0111 21.78
ResShift-15 [yue2024resshift] 35.84 0.9828 0.0097 6.54
QDM-S-15 33.99 0.9766 0.0162 0.74
QDM-B-15 35.10 0.9804 0.0147 2.84
QDM-L-15 35.75 0.9824 0.0120 10.00
Med-SR8 (64→512) LDM-50 [rombach2022high] 31.00 0.9635 0.0162 21.78
ResShift-15 [yue2024resshift] 32.63 0.9728 0.0156 6.54
QDM-S-15 30.39 0.9617 0.0240 0.68
QDM-B-15 31.69 0.9681 0.0198 2.58
QDM-L-15 33.05 0.9743 0.0159 9.08
@@ 4.2 Model Analysis
We benchmark our method against contemporary super-resolution approaches spanning two dominant paradigms: (1) Pretrained-Prior-Based Methods, including StableSR [wang2024exploiting], DiffBIR [lin2024diffbir], SeeSR [wu2024seesr], OSEDiff [wu2025one], InvSR [yue2024arbitrary] and SinSR [wang2024sinsr], which inherit generative priors from Stable Diffusion [rombach2022high, podell2023sdxl] or employ distillation from a pretrained model; (2) Training-from-scratch Methods, comprising BSRGAN [zhang2021designing], RealESRGAN [wang2021real], and SwinIR [liang2021swinir], LDM [rombach2022high], and ResShift [yue2024resshift]. For sampling consistency, we configure 50 diffusion steps for StableSR, DiffBIR, SeeSR, and LDM. Method-specific configurations are preserved for OSEDiff (1 steps), InvSR (1 steps), SinSR (1 step), and ResShift (15 steps) following their official implementations.
[…]
Dataset Metric LDM ResShift QDM-L
[…]
SegRap CECT
(Out-of-Distribution)
[…]
As demonstrated by the qualitative and quantitative results in Fig. 4, Table 1, and Table 2, our proposed model outperforms or achieves comparable performance to state-of-the-art training-from-scratch methods across multiple benchmarks. While it still falls short of prior-based methods, it is worth emphasizing that such methods face significant practical limitations in medical imaging applications due to the absence of foundational generative models (e.g., Stable Diffusion) tailored to this domain. Furthermore, Table 2 highlights the computational efficiency of our QDM-B/S model: it delivers competitive performance relative to LDM and ResShift while requiring substantially less computational overhead. In contrast, lightweight methods like BSRGAN and SwinIR, though computationally efficient, achieve markedly inferior results. More importantly, Table 3 demonstrates that QDM attains superior PSNR and SSIM scores in tumor regions on both SegRap Tumor CTs and the more challenging out-of-distribution SegRap Tumor CECTs, outperforming LDM and ResShift. This underscores QDM’s effectiveness in enhancing clinically critical pathological areas. Additional qualitative results are provided in Appendix D.
@@ 4.3 Ablation Studies
Quadtree mask threshold. As detailed in Section 3.3, the quadtree mask density controls computation allocation by adaptively partitioning images into homogeneous regions (processed via efficient upstream) and detail-rich areas (handled by high-resolution downstream). Table 4 demonstrates this on Med-SR8: At $s=0$ (lossless mode), QDM-L matches full-mask performance while cutting computation by $40.3\%$, confirming that non-uniform prediction suffices for medical SR. Increasing $s$ to $0.90$ reduces computation to $15.5\%$ baseline ($6.5\times$ less) with only $0.34$ dB PSNR drop, favoring efficiency-critical clinical use. To balance performance and efficiency, we set the threshold $s=0.15$ for all medical SR experiments.
@@
Threshold PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ %Density FLOPs(T)
Full Mask 33.06 0.9744 0.0144 100.00 36.64
0.00 33.06 0.9744 0.0144 49.14 21.88
0.15 33.05 0.9743 0.0159 9.59 9.08
0.30 33.01 0.9739 0.0174 7.64 8.18
0.60 32.85 0.9728 0.0201 4.56 7.30
0.90 32.67 0.9719 0.0221 2.25 5.68
@@
Benchmark Param PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FLOPs(T)
Med-SR4 (128→512) Patch (4) 35.18 0.9806 0.0113 8.20
Patch (8) 35.10 0.9804 0.0115 6.28
Window (4) 34.91 0.9797 0.0119 5.98
Window (8) 35.10 0.9804 0.0115 6.28
Window (16) 35.15 0.9807 0.0114 7.16
Med-SR8 (64→512) Patch (4) 31.88 0.9690 0.0172 8.14
Patch (8) 31.69 0.9682 0.0178 6.22
Window (4) 31.55 0.9673 0.0181 5.88
Window (8) 31.69 0.9682 0.0178 6.22
Window (16) 31.73 0.9683 0.0175 7.12
@@ 5 Conclusion
In this work, we present QDM, a novel diffusion-based super-resolution (SR) method that leverages quadtree-based region-adaptive sparse diffusion to enhance detail-rich image regions while minimizing computational overhead in homogeneous areas. The proposed framework incorporates two key innovations: 1) a quadtree-based sparse diffusion model that rapidly detects detail-rich regions and only performs diffusion process in such areas, and 2) a two-stream architecture that decouples global and local feature extraction. The upstream branch captures broad contextual information using large patch sizes, while the downstream branch employs cross-attention to refine high-detail chunks identified via quadtree masks. This design achieves two advantages. First, it eliminates the need to store high-dimensional feature maps, significantly reducing memory overhead. Second, it dynamically minimizes computation by selectively processing only salient regions in the downstream branch. Extensive experiments on real-world SR and medical CT reconstruction tasks demonstrate that QDM outperforms existing training-from-scratch diffusion models in both reconstruction quality and computational efficiency. These results underscore the potential of region-adaptive diffusion strategies for resource-conscious high-fidelity image enhancement.
@@ Appendix A Different Configurations of QDM
Table 6: Details of QDM models. Our configurations for Small (S), Base (B) and Large (L) models are based on those of the DiT model [peebles2023scalable]. To achieve a comparable parameter count, we halve the number of layers from the original DiT configurations.
Model Layers NN Hidden size Heads #Parameters(M)
@@ Appendix B Implementation Details
Data Processing. For the real-world SR task, we adopt fixed-size $512\!\times\!512$ image crops during training. Following recent work [yang2024inf], training images exceeding this resolution undergo either: (1) direct random cropping, or (2) resizing the shorter edge to 512 pixels before random cropping. We synthesize LR/HR pairs using the degradation pipeline from Real-ESRGAN [wang2021real].
Beyond real-world test benchmarks, we construct LSDIR-Test using center-cropped $512\!\times\!512$ images from the LSDIR test set [li2023lsdir], processed through identical degradation parameters as the training data.
For the medical CT SR task, we use the first degradation stage of Real-ESRGAN pipeline [wang2021real] simulating clinical noise artifacts. Both training and test images are first downsampled to $512\!\times\!512$ resolution as HR references, then processed through the adopted pipeline to obtain the corresponding LR image. Thus, we can create two benchmarks: Med-SR4 ($\times 4$ super-resolution task) and Med-SR8 ($\times 8$ super-resolution task).
Evaluation Metrics. The performance of different methods was evaluated using both reference-based and non-reference metrics. For reference-based assessment, we employed PSNR, SSIM [wang2004image], and LPIPS [zhang2018unreasonable]. To better align with human perceptual judgments in generative super-resolution, we also incorporated non-reference metrics: CLIPIQA [wang2023exploring] and MUSIQ [ke2021musiq]. All LPIPS, CLIPIQA, and MUSIQ results were computed following the official implementations provided in IQA-PyTorch [pyiqa] [^1] https://github.com/chaofengc/IQA-PyTorch.
Training Process & Hyperparamters. In real-world SR task, following prior work (LDM [rombach2022high], ResShift [yue2024resshift]), our architecture operates in latent space using a Vector Quantized GAN (VQGAN, [esser2021taming]) with a downsampling factor of 4. The model was trained for 150,000 iterations using the dual-stream training objective (Eq. 6), with a batch size of 64 on eight NVIDIA A100 80GB PCIe GPUs. We employed the Adam optimizer with a learning rate of $5\times 10^{-5}$, followed by a 50,000-iteration fine-tuning phase using the perceptual training objective (Eq. 7). For all real-world experiments, we set the quadtree threshold $s=0.00$.
[…]
Besides, Figure 6 provides a detailed structure of cross-attention block presented in Figure 2. Each cross-attention block processes features sequentially through three components: a local intra-window self-attention, a cross-attention mechanism that integrates global context from upstream features, and a feedforward network. It’s important to note that the upstream global features are pre-projected into reusable key-value (kv) pairs to minimize redundant computations. The MLP ratio is set to the default value of 4.
Ultra-High-Resolution Inference. To handle ultra-high-resolution (UHR) images beyond GPU memory limits, we adopt a patch-based inference strategy with Gaussian-weighted fusion proposed in [yue2024arbitrary]. The input image $I_{\mathrm{LR}}\in\mathbb{R}^{B\times C\times H\times W}$ is divided into overlapping patches of size $P\times P$ and stride $S\leq P$. Each patch is super-resolved individually to $\hat{p}_{k}\in\mathbb{R}^{C\times(s_{f}P)\times(s_{f}P)}$ and accumulated into a global canvas with corresponding Gaussian weights $w_{k}(x)$ generated by separable 1D kernels. The final output is obtained by normalized fusion:
$I_{\mathrm{HR}}(x)=\frac{\sum_{k}w_{k}(x)\,\hat{p}_{k}(x)}{\sum_{k}w_{k}(x)}.$
This Gaussian overlap-add formulation ensures seamless blending between neighboring patches while enabling scalable processing of arbitrarily large images. In practice, we set P=128P\!=128, S=112S\!=\!112, and accumulate in float64 precision for numerical stability.